# Spectral representations of interpolation spaces of reproducing kernel Hilbert spaces

### v1

@@ A Appendix: Interpolation spaces
We begin by proving that isometric isomorphisms can be carried over from Banach spaces $E_{2}\hookrightarrow E_{1}$ to the interpolation spaces $E_{1},E_{2}]_{\theta,r}$. For a map $f:A\rightarrow B$ and $A^{\prime}\subset A$, $B^{\prime}\subset B$, we use the notation $f|_{A^{\prime}\rightarrow B^{\prime}}$ for the map $A^{\prime}\rightarrow B^{\prime}$ given by $x\mapsto f(x)$ in case it is well defined.
@@
Let $E_{2}\hookrightarrow E_{1}$ and $F_{2}\hookrightarrow F_{1}$ be Banach spaces. Suppose there is a bounded linear operator $\phi:E_{1}\rightarrow F_{1}$ such that $\phi|_{E_{1}\rightarrow F_{1}}$ and $\phi|_{E_{2}\rightarrow F_{2}}$ are isometric isomorphisms. Then for $0<\theta<1$ and $1\leq r\leq\infty$, the following map is an isometric isomorphism:
@@
$[H,H_{2}]_{\theta,r}\cup[H_{1},H_{2}]_{\theta,r}\subset H_{1}$ (120)
@@ B Appendix: Direct sums of Banach spaces
A closer look at Definition 2.1 reveals that the $\ell^{(\mu,\theta,r)}$-norms are built up from two nested norms: The inner part (10) of the $(\mu,\theta,r)$-norm originates from an $\ell^{2}(\mu^{-\theta})$-norm taken over the restricted indices in $M_{j}$, while the outer sum in (11) corresponds to the $\ell^{r}$-norm taken over the values given by the inner part.
@@
Thereby, $\Leftarrow$ and $\leq$ follow directly. To see $\Rightarrow$ and $\geq$, we choose $\tilde{x}_{i}:=x_{i}$ if $x_{i}^{\prime}(x_{i})\geq 0$ and $\tilde{x}_{i}:=-x_{i}$ otherwise. Then we find that $x_{i}^{\prime}(\tilde{x}_{i})=|x_{i}^{\prime}(x_{i})|$, and we still have $\|\tilde{x}\|_{p}=\|x\|_{p}$. Hence using $\tilde{x}$ instead of $x$ on the left hand side gives the right hand side.
@@
    isometric isomorphism if $p=0$ and $p^{\prime}=1$.
@@
    isometric isomorphism if $1\leq p<\infty$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
@@
    isometry if $p=\infty$ and $p^{\prime}=1$.
@@
A proof of (i) and (ii) can be found in [1, Theorem 16.49 and the following remark]. Let us prove (iii) here. We begin by showing that the operator is well defined in the sense that the sum converges. In fact, we even have absolute convergence, as for $x^{\prime}\in{\bigoplus_{i\in I}}^{(1)}E_{i}^{\prime}$ and $x\in{\bigoplus_{i\in I}}^{(\infty)}E_{i}$, we find that
@@
where we used the absolute convergence of (145). Hence the uniform boundedness principle [13, Theorem 4.52] yields that $\sup_{n\in\mathbb{N}}\|S_{n}\|<\infty$.
@@
$\|S_{n}\|=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}((x_{i}\mathbbm{1}_{i\leq n})_{i\in\mathbb{N}}|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\leq n}^{(0)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}\text{.}$ (150)
This means that the operator norm of $S_{n}$ equals the operator norm of $S_{n}$ restricted to ${\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i}$. This restricted operator is as an element of $({\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i})^{\prime}\simeq{\bigoplus_{i\in\mathbb{N}}}^{(1)}E_{i}^{\prime}$, where we are using the isometric isomorphism from Lemma B.3. This isometric isomorphism assigns $S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}$ to the sequence $(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\in\bigoplus_{i\leq n}^{(1)}E_{i}^{\prime}$. We hence find the equality of norms
$\|S_{n}\|=\left\|S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}\right\|=\left\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\right\|_{1}\text{.}$ (151)
@@
    It holds that $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$ with norm $\|x^{\prime}\|_{p^{\prime}}\leq C$.

### v2

@@ A Appendix: Interpolation spaces
We begin by proving that isometric isomorphisms can be carried over from Banach spaces $E_{2}\hookrightarrow E_{1}$ to the interpolation spaces $E_{1},E_{2}]_{\theta,r}$. For a map $f:A\rightarrow B$ and $A^{\prime}\subset A$, $B^{\prime}\subset B$, we use the notation $f|_{A^{\prime}\rightarrow B^{\prime}}$ for the map $A^{\prime}\rightarrow B^{\prime}$ given by $x\mapsto f(x)$ in case it is well-defined.
@@
Let $E_{2}\hookrightarrow E_{1}$ and $F_{2}\hookrightarrow F_{1}$ be Banach spaces. Suppose there is a bounded linear operator $\phi:E_{1}\rightarrow F_{1}$ such that $\phi$ and $\phi|_{E_{2}\rightarrow F_{2}}$ are isometric isomorphisms. Then for $0<\theta<1$ and $1\leq r\leq\infty$, the following map is an isometric isomorphism:
@@
$[H,H_{2}]_{\theta,r}\cup[H_{1},H_{2}]_{\theta,r}\subset H_{1}\text{.}$ (142)
@@ B Appendix: Direct sums of Banach spaces
A closer look at Definition 2.2 reveals that the $\ell^{(\mu,\theta,r)}$-norms are built up from two nested norms: The inner part (10) of the $(\mu,\theta,r)$-norm originates from an $\ell^{2}(\mu^{-\theta})$-norm taken over the restricted indices in $M_{j}$, while the outer part in (11) corresponds to the $\ell^{r}$-norm taken over the values given by the inner part.
@@
Thereby, $\Leftarrow$ and $\leq$ follow directly. To see $\Rightarrow$ and $\geq$, we choose $\tilde{x}_{i}:=x_{i}$ if $x_{i}^{\prime}(x_{i})\geq 0$ and $\tilde{x}_{i}:=-x_{i}$ otherwise. Then we find that $x_{i}^{\prime}(\tilde{x}_{i})=|x_{i}^{\prime}(x_{i})|$, and we still have $\|\tilde{x}\|_{p}=\|x\|_{p}$. Hence using $\tilde{x}$ instead of $x$ on the left-hand side gives the right-hand side.
@@
    an isometric isomorphism if $p=0$ and $p^{\prime}=1$,
@@
    an isometric isomorphism if $1\leq p<\infty$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$, and
@@
    an isometry if $p=\infty$ and $p^{\prime}=1$.
@@
A proof of (i) and (ii) can be found in [1, Theorem 16.49 and the following remark]. Let us prove (iii) here. We begin by showing that the operator is well-defined in the sense that the sum converges. In fact, we even have absolute convergence, as for $x^{\prime}\in{\bigoplus_{i\in I}}^{(1)}E_{i}^{\prime}$ and $x\in{\bigoplus_{i\in I}}^{(\infty)}E_{i}$, we find that
@@
where we used the absolute convergence of (167). Hence the uniform boundedness principle [21, Theorem 4.52] yields that $\sup_{n\in\mathbb{N}}\|S_{n}\|<\infty$.
@@
$\|S_{n}\|=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}((x_{i}\mathbbm{1}_{i\leq n})_{i\in\mathbb{N}}|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\leq n}^{(0)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}$ (172)
This means that the operator norm of $S_{n}$ equals the operator norm of $S_{n}$ restricted to ${\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i}$. This restricted operator is an element of $({\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i})^{\prime}\simeq{\bigoplus_{i\in\mathbb{N}}}^{(1)}E_{i}^{\prime}$, where we are using the isometric isomorphism from Lemma B.3. This isometric isomorphism assigns $S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}$ to the sequence $(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\in\bigoplus_{i\leq n}^{(1)}E_{i}^{\prime}$. We hence find the equality of norms
$\|S_{n}\|=\left\|S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}\right\|=\left\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\right\|_{1}$ (173)
@@
    It holds that $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$ with $\|x^{\prime}\|_{p^{\prime}}\leq C$.
@@
C Appendix: Semicontinuous functions
Let $(X,\mathcal{T})$ be a topological space. Then a function $g:X\rightarrow[0,\infty]$ is called lower semicontinuous (l.s.c.) if and only if the sets $g^{-1}((a,\infty])$ are open in $X$ for all $a\geq 0$.
Lemma C.1.
Let $(X,\mathcal{T})$ be a topological space and suppose that $g_{i}:X\rightarrow[0,\infty]$ are l.s.c. functions for all $i$ in an index set $I$. Then the following functions $X\rightarrow[0,\infty]$ are l.s.c.:
  1. (i)
    $\sup_{i\in I}g_{i}$
  2. (ii)
    $\sum_{i\in I}g_{i}$
  3. (iii)
    $g_{1}\cdot g_{2}$
  4. (iv)
    $g_{1}^{\alpha}$ for all $\alpha>0$
Proof.
(i). Non-negativity follows directly, for lower semicontinuity, see [5, §6.2 Proposition 2].
(ii). Non-negativity follows directly. By [5, §6.2 Proposition 2], finite sums of l.s.c. functions are l.s.c. The property follows for infinite sums as $\sum_{i\in I}g_{i}(x)=\sup_{j\in I}\sum_{i=1}^{j}g_{i}(x)$ holds for non-negative functions.
(iii). Non-negativity follows directly, for lower semicontinuity, see [5, §6.2 Proposition 2].
(iv). This follows from the fact that for all $a\geq 0$ we have
$\{x\in X\,|\,g_{1}(x)^{\alpha}>a\}=\{x\in X\,|\,g_{1}(x)>a^{\frac{1}{\alpha}}\}\text{,}$
which is an open set. ∎
Lemma C.2.
Let $(X,\mathcal{T})$ be a topological space and $(X,\mathcal{A},\nu)$ be a measure space such that $\nu$ is $\mathcal{T}$-positive. Then for all l.s.c. functions $g:X\rightarrow[0,\infty]$ it holds that
$\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)=\sup_{x\in X}g(x)\text{.}$ (181)
Proof.
Assume the converse, that means $\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)<\sup_{x\in X}g(x)$. It follows that there are $\varepsilon>0$ and $y\in X$ such that the set
$M:=\{y\in X\,|\,g(y)-\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)\in(\varepsilon,\infty]\}$ (182)
is not empty. Additionally, $M$ is open as $g$ is l.s.c. By the positivity of $\nu$, we find that $\nu(M)>0$. Hence $g>\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)$ holds on the set of positive measure $M$, a contradiction. ∎
Corollary C.3.
Let $(X,\mathcal{T})$ be a topological space and $(X,\mathcal{A},\nu)$ be a measure space such that $\nu$ is $\mathcal{T}$-positive. Let $f:X\rightarrow\mathbb{R}$ be continuous. Then $\|f\|_{\mathcal{L}^{\infty}(X)}=\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}$.
Proof.
Apply Lemma C.2 to the non-negative and $\mathcal{T}$-continuous function $|f|$. ∎
Lemma C.4.
Let $(X,\mathcal{T})$ be a topological space and $(X,\mathcal{A},\nu)$ be a measure space such that $\nu$ is $\mathcal{T}$-positive. Then $\{[f]_{\sim}\,|\,f\in\mathrm{C}_{b}(X,\mathcal{T})\}$ is a closed subspace of $\mathrm{L}^{\infty}(\nu)$ which is isometric isomorphic to $\mathrm{C}_{b}(X,\mathcal{T})$ by the mapping $f\mapsto[f]_{\sim}$ for $f\in\mathrm{C}_{b}(X,\mathcal{T})$.
Proof.
The fact that $\{[f]_{\sim}\,|\,f\in\mathrm{C}_{b}(X,\mathcal{T})\}$ is a linear subspace of $\mathrm{L}^{\infty}(\nu)$ follows from the definition. It remains to prove that it is closed. We do so by proving completeness, which is equivalent to closedness. To this end, let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of functions in $\mathrm{C}_{b}(X,\mathcal{T})$ such that the $f_{n}]_{\sim}$ form a Cauchy sequence in $\mathrm{L}^{\infty}(\nu)$-norm. It follows by Corollary C.3 that for $n,m\rightarrow\infty$, we have
$\|f_{n}-f_{m}\|_{\mathcal{L}^{\infty}(X)}=\|[f_{n}-f_{m}]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|[f_{n}]_{\sim}-[f_{m}]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}\rightarrow 0\text{.}$ (183)
As $\mathrm{C}_{b}(X,\mathcal{T})$ with the $\mathcal{L}^{\infty}(X)$-norm is complete, there is a $f\in\mathrm{C}_{b}(X,\mathcal{T})$ such that $\|f_{n}-f\|_{\mathcal{L}^{\infty}(X)}\rightarrow 0$. But this implies
$\|[f_{n}]_{\sim}-[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|[f_{n}-f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|f_{n}-f\|_{\mathcal{L}^{\infty}(X)}\rightarrow 0$ (184)
by another application of Corollary C.3. This shows that $f]_{\sim}$ is the desired limit of $([f_{n}]_{\sim})_{n\in\mathbb{N}}$ in $\mathrm{L}^{\infty}(\nu)$-norm.
Finally, as the $\mathcal{L}^{\infty}(X)$-norm of a function $f\in\mathrm{C}_{b}(X,\tau(H))$ agrees with the $\mathrm{L}^{\infty}(\nu)$-norm of $f]_{\sim}$ by Corollary C.3, the map
$\mathrm{C}_{b}(X,\mathcal{T})\rightarrow\{[f]_{\sim}\,|\,f\in\mathrm{C}_{b}(X,\mathcal{T})\},\quad f\mapsto[f]_{\sim}$ (185)
is isometric, hence injective. Its surjectivity follows directly. ∎